OpenStudy (anonymous):

Solve using law of cosines or law of sines A=50* B=30* c=9* a= b= C=

OpenStudy (anonymous):

@myininaya when you have time

myininaya (myininaya):

You are given ASA that means you are given a side between 2 angles law of sines

myininaya (myininaya):

http://home.windstream.net/okrebs/page94.html here is cool site to see which goes with which

OpenStudy (anonymous):

Okay i tried to figure it out but I cant

myininaya (myininaya):

find that 3rd angle first

myininaya (myininaya):

I was start with finding the 3rd angle if given the first 2

myininaya (myininaya):

once you find the 3rd angle use the law of sines

OpenStudy (anonymous):

so c^2 =b^2-a^2?

myininaya (myininaya):

180=A+B+C

OpenStudy (anonymous):

im sorry +

OpenStudy (anonymous):

Wait, huh ?

myininaya (myininaya):

Do you know the sum of the angles in a triangle is 180 degrees?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

180=50+30+C?

myininaya (myininaya):

So you know A+B+C=180 right

OpenStudy (anonymous):

80=c?

OpenStudy (anonymous):

100 ****

myininaya (myininaya):

100 is C not c

OpenStudy (anonymous):

C=100

myininaya (myininaya):

Right so now you have all 3 angles and you also have side c

myininaya (myininaya):

so now you can go into using law of sines

myininaya (myininaya):

you can do either \[\frac{\sin(A)}{a}=\frac{\sin(C)}{c} \text{ or } \frac{\sin(B)}{b}=\frac{\sin(C)}{c}\]

myininaya (myininaya):

actually go ahead and do both side we need both a and b

OpenStudy (anonymous):

Can i just type it in my calculator?

OpenStudy (anonymous):

Crap nevermind we dont have a or b

myininaya (myininaya):

first solve the first one for a and solve the second one for b then you can use the calculator

OpenStudy (anonymous):

\[\frac{ \sin 50 }{ a }=\frac{ \sin 100 }{ 9 }\] ?

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

yes now cross multiply

OpenStudy (anonymous):

9sin50=asin100

myininaya (myininaya):

ok and the divide by sin(100)

myininaya (myininaya):

you are ready to put into your calculator now

OpenStudy (anonymous):

7.0007?

myininaya (myininaya):

yah

OpenStudy (anonymous):

Okay what next :)

myininaya (myininaya):

find b

myininaya (myininaya):

\[\frac{\sin(B)}{b}=\frac{\sin(C)}{c} \]

OpenStudy (anonymous):

4.57=b

myininaya (myininaya):

sounds great

myininaya (myininaya):

that's it

myininaya (myininaya):

good job

OpenStudy (anonymous):

Thank you.

myininaya (myininaya):

np